there are three bags first containig 1 white 2 red 3 green ,second 2 white 3 red 1 green balls and third 3 white 1 red 2 green balls. two balls drawn from a bag chosen at random. these are found to be one red and one white.find the probability that the balls so drawn came from second bag.

Grade:Upto college level

1 Answers

Chetan Mandayam Nayakar
13 years ago

answer:(2/6)(3/6)/((2/6)(3/6)+(1/6)(3/6)+(1/6)(2/6)) =6/(6+2+3)=6/11